OpenStudy (anonymous):

I really need the help Match each actual value with the predicted value that results in a thirty−five percent error. Actual: one hundred forty A) Predicted:ninety−one Actual: one hundred B) Predicted:seventy−eight Actual: eighty C) Predicted: sixty−five Actual: one hundred twenty D) Predicted: one hundred eight

OpenStudy (anonymous):

Basically, a 35% error means that the actual value may be either 35% higher or lower than the latter. For example, 20% of X is \[\frac{ 20 }{ 100 } \times x = 0.2x\] Therefore, the margin of error will be \[x-0.2x \le x \le x + 0.2x\] Let's say x = 200 and the 20% error still the same Therefore, the range of errors lie between: \[200 - 0.2(200) \le 200 \le 200 + 0.2(200) \] I hope that helps

OpenStudy (anonymous):

What I meant is that the predicted value will be between the margin of 35% higher or lower than the actual value.